OpenStudy (anonymous):

Factor completely: 8x^2+20x-12

OpenStudy (anonymous):

(8x-4)(x+3) :)

OpenStudy (anonymous):

thanks. can you factorise 2t^2-18-20

OpenStudy (anonymous):

(2t-20)(t+1)

OpenStudy (anonymous):

thank you soooo much. can you possibly factorise 8t^2+20x-12 ;)

OpenStudy (anonymous):

\[\frac{ x ^{2}+5x+4 }{ x ^{2}+x-12 }\]

OpenStudy (anonymous):

Divide all by 4. 2x^2+5x-3 =(2x-1)(x+3)

OpenStudy (anonymous):

\[\frac{ \left( x+1 \right) \left( x+4 \right)}{\left( x-3 \right)\left( x+4 \right)}\] \[= \frac{ \left( x+1 \right)}{ \left( x-3 \right) }\]